# Human Insights Driven Latent Space for Different Driving Perspectives: A Unified Encoder for Efficient Multi-Task Inference

### v1

@@ Abstract
Autonomous driving holds great potential to transform road safety and traffic efficiency by minimizing human error and reducing congestion. A key challenge in realizing this potential is the accurate estimation of steering angles, which is essential for effective vehicle navigation and control. Recent breakthroughs in deep learning have made it possible to estimate steering angles directly from raw camera inputs. However, the limited available navigation data can hinder optimal feature learning, impacting the system’s performance in complex driving scenarios. In this paper, we propose a shared encoder trained on multiple computer vision tasks critical for urban navigation, such as depth, pose, and 3D scene flow estimation, as well as semantic, instance, panoptic, and motion segmentation. By incorporating diverse visual information used by humans during navigation, this unified encoder might enhance steering angle estimation. To achieve effective multi-task learning within a single encoder, we introduce a multi-scale feature network for pose estimation to improve depth learning. Additionally, we employ knowledge distillation from a multi-backbone model pretrained on these navigation tasks to stabilize training and boost performance. Our findings demonstrate that a shared backbone trained on diverse visual tasks is capable of providing overall perception capabilities. While our performance in steering angle estimation is comparable to existing methods, the integration of human-like perception through multi-task learning holds significant potential for advancing autonomous driving systems. More details and the pretrained model are available at https://hi-computervision.github.io/uni-encoder/.
@@ I Introduction
The development of autonomous driving capabilities requires comprehensive understanding, reliable real-time decision-making, and robust control. Traditional approaches to tasks like steering control rely on modular architectures that separate perception, decision-making, and control, each requiring domain-specific expertise and extensive calibration to function effectively [1]. However, recent advancements in deep learning have introduced end-to-end systems, where neural networks directly learn steering commands from raw input data, greatly simplifying the design and reducing the need for manual tuning [2, 3, 4, 5].
Most deep learning-based steering estimation uses Convolutional Neural Networks (CNN) to extract features from RGB images [6, 7, 8, 9, 10]. While these methods demonstrate that CNNs can extract useful information for estimating the steering angle, the training datasets may lack sufficient task-specific details for comprehensive learning. Indeed, Capito et al. [11] demonstrate that using optical flow from the same RGB images as an additional input can significantly enhance the original method. This finding raises questions about the scene information encoded by the encoder and whether it effectively captures human-like visual perception needed for navigation.
In this paper, we propose a new approach that integrates human-like perceptual information into a single encoder through a multi-task training strategy. The computer vision tasks we focus on are crucial for urban navigation, including depth, pose and 3D scene flow estimation, as well as semantic, instance, panoptic, and motion segmentation (see Figure 1). Our method offers several key advantages. First, after the initial training, the model can generate multiple outputs with a single pass through the shared encoder, reducing inference time. Additionally, our experimental results demonstrate that this approach yields competitive performance compared to state-of-the-art methods for each individual task. Finally, the encoder model can be frozen and stacked with a prediction head for steering estimation, allowing us to evaluate the relevance of the learned features for this navigation task.
@@
  • A novel pose decoder architecture that leverages the multi-scale features from our shared encoder to enhance depth estimation.
  • A strategy to stabilize training by distilling knowledge from a multi-encoder teacher pretrained on same tasks.
  • An evaluation of whether a shared encoder trained on multiple tasks can improve steering angle estimation.
@@ II-B Monocular Depth & Pose Estimation
Monocular depth estimation is a task of predicting the depth of a scene from a single 2D image. Unlike stereoscopic methods, which use different viewpoints as input, monocular depth estimation must infer depth from just one, making it particularly challenging. Traditional methods mostly based on hand-crafted features, which can lead to inaccuracies in complex senarios [15]. In a first attempt using deep learning, Eigen et al. (2014) introduced a CNN-based model to predict depth maps directly from single images, achieving superior results. However, this method requires ground truth for depth, typically obtained with expensive hardware like LiDAR, limiting its practicality. To overcome this, recent research has shifted to unsupervised methods that use the inherent structure of unlabeled images. Notably, Zhou et al. [16] and Godard et al. [17] proposed unsupervised methods which employ video sequences to simultaneously learn depth and camera motion, reducing the dependency on labeled data and adapting better to dynamic scenes.
However, unsupervised methods often consider the assumption of a static world, which is not true in most real senarios. To this end, recent advancements have intergrated flow (e.g., 2D or 3D) and motion segmentation in addition to depth and pose estimation, enabling better handling of dynamic objects in the scene [18].
@@ II-C Scene Flow and Motion Segmentation
II-D Steering angle estimation
The two main research directions for steering estimation are model-based and model-free. Model-based methods rely on vehicle dynamics models, while model-free methods leverage data-driven techniques like deep learning. This paper focuses only on the model-free approach. Bojarski et al. [8] introduced an end-to-end steering angle estimation using CNN with a single RGB camera input. Capito et al. [11] improved the navigation capability by incorporating optical flow into the input. The importance of temporal information has also been highlighted in many works. Indeed, Eraqi et al. [22] employed LSTM networks to enhance steering control, while Xu et al. [23] combined fully convolutional networks with LSTM and semantic segmentation for improved road condition interpretation. Lechner et al. [2] proposed Neural Circuit Policies (NCPs), which provide interpretable decision maps from high-dimensional inputs. These studies indicate that integrating temporal information between frames can improve steering accuracy.
@@ III Methods
Due to the lack of a comprehensive dataset with annotations for all tasks, we used the CityScapes dataset [25], which provides high-quality segmentation annotations and a video sequence format suitable for tasks like depth and pose estimation, 3D scene flow, and motion mask segmentation.
III-A Panoptic, Instance, and Semantic Segmentations
For these three supervised segmentation tasks, we use the OneFormer approach. This is a task-conditioned universal image segmentation model that achieves state-of-the-art performance across all segmentation tasks with a unified model and architecture. The OneFormer framework adopts a task-conditioned joint training strategy that allows simultaneous training on all segmentation tasks using a single model, reducing resource requirements significantly.
@@ III-A Panoptic, Instance, and Semantic Segmentations
III-B Depth, Pose, 3D Flow Estimation, and Motion Mask
@@ IV-A Training Setup
We use the KITTI Eigen split for initial ablation and CityScapes for the remaining experiments, with image resolutions of $192\times 640$ and $192\times 512$, respectively. Images are preprocessed into triples using scripts from [16]. For CityScapes, the lower 25% of images is cropped to exclude the front car [28]. The entire CityScapes dataset is used for supervised segmentation tasks. Pretraining is conducted on a single NVIDIA RTX A5000 with a batch size of 6 (3 images for segmentation and 3 triples for other tasks) over four steps: (1) training the shared encoder with depth, pose, and segmentation decoders for 60,000 steps, (2) training the 3D scene flow decoder for 40,000 steps with other components frozen, (3) training the pose decoder, 3D scene flow decoder, and motion mask segmentation decoder for 40,000 steps with other components frozen, and (4) training the whole network for 250,000 steps. After pretraining, we follow the tenfold cross-validation procedure by Lechner et al. [2] to fine-tune the model on the steering angle estimation task. The values of hyperparameters are: $\lambda_{\text{cls}}=2,\lambda_{\text{bce}}=\lambda_{\text{dice}}=\lambda_{\text{c}}=5,\lambda_{\text{contrast}}=0.5,\lambda_{\text{m}}=\lambda_{\text{g}}=\beta_{1}=0.1,\beta_{2}=1e-3,\lambda_{\text{1}}=\lambda_{\text{2}}=\lambda_{\text{3}}=1$.
@@ IV-B Ablation Studies
1 $MD2_{separate}$ 0.115 0.193 0.877
$MD2_{shared}$ 0.125 0.201 0.857
$\text{Our}_{separate}$ 0.108 0.183 0.884
$\text{Our}_{shared}$ 0.117 0.190 0.872
$\text{Our}_{shared}$ 0.109 0.184 0.884
2 $\text{Our}_{separate}$ 0.103 0.157 0.885
$\text{Our}_{shared}$ 0.126 0.182 0.850
$\text{Our}_{shared}$ 0.134 0.183 0.833
$\text{Our}_{shared}$ 0.106 0.158 0.888
TABLE I: 1) Ablation on multi-scale pose decoder using KITTI dataset. 2) Analysis on knowledge distilation on CityScapes. MS, FM, and KD refers multi-scale pose decoder, 3D scene Flow & Motion mask, and Knowledge Distilation respectively. For the separate encoder, a ResNet18 was employed similar to [17].
@@
Method IM Semantics #$\#f$ D Error metric ($\downarrow$) Accuracy metric ($\uparrow$)
at inference $Abs_{rel}$ $Sq_{rel}$ $RMSE$ $RMSE_{log}$ $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
@@
TABLE III: Depth evaluation on the KITTI (K), and CityScapes (CS) Dataset. IM stands for independent motion. #$\#f$ indicates the number of frames during inference. D is the used datatset for training and evaluation. Bold is best. Underline is ours.
Knowledge distillation for 3D flow & motion mask. We evaluate the effectiveness of knowledge distillation in training our shared encoder on the CityScapes dataset, which includes many dynamic objects, making 3D scene flow and motion mask networks crucial. Table I-2 shows the depth error and accuracy for different models. First, using a shared encoder with our multi-scale pose decoder achieves good performance compared to using separate encoders for different tasks. Second, training the shared encoder without knowledge distillation leads to suboptimal performance, even worse than the shared encoder model that doesn’t consider 3D scene flow and motion masks. However, incorporating knowledge distillation improves the shared encoder’s performance, making its performance inline with models with separate encoders. This highlights the effectiveness of our strategy.
@@ IV-C Comparison with State-of-the-Art Methods
CNN-GRU (64 units) $1.25\pm 1.02$ $5.06\pm 6.64$
CNN-LSTM (64 units) $\mathbb{0.19}\pm 0.05$ $\mathbb{3.17}\pm 3.85$
VAE-LSTM (64 units) $0.54\pm 0.26$ $4.70\pm 4.80$
VAE-LSTM (19 units) $0.60\pm 0.30$ $6.75\pm 8.33$
Swin-AttnPool (Ours) $3.20\pm 1.50$ $6.15\pm 4.86$
@@
IV-D From Dense Latent Space to Steering Angle: Evaluation
Table IV showcases the performance of various models for steering prediction, including our proposed Swin-AttnPool approach and several existing CNN-based [2] and VAE-based models [5]. While our model shows higher training and test errors compared to some of the baseline models, it is important to note that our approach is training the decision module with direct latent space from frozen encoder and aims to leverage a more comprehensive and human-like understanding of the real driving environment.
V Discussion and Conclusion
Training all tasks within a single encoder using a naive approach often leads to sub-optimal performance for some tasks. Our method is one of the first attempts to incorporate more human-like perception information into a shared encoder, achieving performance comparable to multi-encoder methods in the literature.
We argue that accurate driving requires human-like perception and decision-making. To enhance the decision-making model, we propose using a dense feature space enriched with human-understandable features or those that can infer such features. While this approach may show higher numerical errors in certain evaluations, it could offer a richer and more contextual understanding of the driving scene, leading to more robust and interpretable decisions in complex real-world scenarios. The higher errors may result from the model learning to consider a broader range of factors beyond just lane-keeping, which could ultimately benefit overall driving performance but is not fully reflected by the current metric.

### v2

@@ Abstract
Autonomous driving systems require a comprehensive understanding of the environment, achieved by extracting visual features essential for perception, planning, and control. However, models trained solely on single-task objectives or generic datasets often lack the contextual information needed for robust performance in complex driving scenarios. In this work, we propose a unified encoder trained on multiple computer vision tasks crucial for urban driving, including depth, pose, and 3D scene flow estimation, as well as semantic, instance, panoptic, and motion segmentation. By integrating these diverse visual cues—similar to human perceptual mechanisms—the encoder captures rich features that enhance navigation-related predictions. We evaluate the model on steering estimation as a downstream task, leveraging its dense latent space. To ensure efficient multi-task learning, we introduce a multi-scale feature network for pose estimation and apply knowledge distillation from a multi-backbone teacher model. Our findings highlight two key findings: (1) the unified encoder achieves competitive performance across all visual perception tasks, demonstrating strong generalization capabilities; and (2) for steering estimation, the frozen unified encoder—leveraging dense latent representations—outperforms both its fine-tuned counterpart and the same frozen model pretrained on generic datasets like ImageNet. These results underline the significance of task-specific visual features and demonstrate the promise of multi-task learning in advancing autonomous driving systems. More details and the pretrained model are available at https://hi-computervision.github.io/uni-encoder/.
@@ I Introduction
The advancement of self-driving cars has gained significant public attention in recent years, though the development of this technology began decades ago. Early innovations include vehicle-to-vehicle communication via radio waves in the 1920s [1] and electromagnetic guidance in the 1930s [2]. The primary goal of autonomous vehicles is to improve road safety and efficiency by minimizing human error, which causes more than 90% accidents in vehicles, while mechanical failures account for only 2% [3]. For fully autonomous driving cars, achieving human-level driving requires comprehensive environmental understanding, robust control, and reliable real-time decision-making.
A fully autonomous driving system usually extract a wide range of visual features—such as depth, motion, and segmentation—to support essential tasks like object detection, path planning, and behavior prediction. Beyond these explicit outputs, intermediate visual features (e.g., extracted by CNNs) that capture spatial relationships, temporal consistency, and scene dynamics are crucial for building a coherent understanding of the environment. These features enable downstream tasks, including obstacle avoidance, adaptive control, and trajectory prediction [4, 5, 6, 7]. Additionally, they play an essential role in interactive human-machine interfaces, which help explain how the system perceives and interprets its surroundings.
However, most learning-based models still focus on single-task objectives, such as steering estimation, using CNNs trained with RGB data to directly minimize the steering error [8, 9, 10, 11, 12]. While these approaches can capture relevant information, their reliance on narrow, task-specific datasets limits their capabilities. For example, Capito et al. [13] demonstrated that incorporating optical flow alongside RGB inputs significantly improved steering performance. This suggests that richer visual cues, beyond simple RGB data, can enhance both steering predictions and the overall contextual understanding of the driving scene.
In this paper, we propose a novel approach that integrates human-like perceptual information into a single encoder using a multi-task training strategy. We focus on computer vision tasks essential for urban navigation, including depth, pose, and 3D scene flow estimation, as well as semantic, instance, panoptic, and motion segmentation (see Figure 1). Our method provides several key advantages. First, after the initial training, the model can generate multiple outputs with a single pass through the unified encoder, reducing inference time while delivering diverse visual outputs. Additionally, our experimental results demonstrate that this approach achieves competitive performance compared to state-of-the-art methods across all individual tasks, ensuring the reliability of the extracted features. Finally, the encoder can be frozen and extended with a prediction head for steering estimation, enabling to assess the relevance of the learned features for this navigation task.
@@
  • A multi-scale pose decoder: We design a pose decoder that leverages multi-scale features from the shared encoder, improving depth estimation in dynamic scenes.
  • Knowledge distillation for stability: We apply knowledge distillation from a multi-encoder teacher model to guide the learning of self-supervised tasks and prevent gradient imbalances.
  • Comprehensive evaluation of shared features: We conduct a comprehensive evaluation of the unified encoder’s performance across all integrated tasks. Our results demonstrate that, for steering angle estimation, the frozen unified encoder outperforms its fine-tuned counterpart and the same architecture pretrained solely on generic datasets such as ImageNet.
@@ II-B Monocular Depth & Pose Estimation
Monocular depth estimation is a task of predicting the depth of a scene from a single 2D image. Unlike stereoscopic methods, which use different viewpoints as input, monocular depth estimation must infer depth from just one, making it particularly challenging. Traditional methods mostly based on hand-crafted features, which can lead to inaccuracies in complex scenarios [17]. In a first attempt using deep learning, Eigen et al. (2014) introduced a CNN-based model to predict depth maps directly from single images, achieving superior results. However, this method requires ground truth for depth, typically obtained with expensive hardware like LiDAR, limiting its practicality. To overcome this, recent research has shifted to unsupervised methods that use the inherent structure of unlabeled images. Notably, Zhou et al. [18] and Godard et al. [19] proposed unsupervised methods which employ video sequences to simultaneously learn depth and camera motion, reducing the dependency on labeled data and adapting better to dynamic scenes.
However, unsupervised methods often consider the assumption of a static world, which is not true in most real scenarios. To this end, recent advancements have intergrated flow (e.g., 2D or 3D) and motion segmentation in addition to depth and pose estimation, enabling better handling of dynamic objects in the scene [20].
@@ II-C Scene Flow and Motion Segmentation
II-D Steering estimation
The two main research directions for steering estimation are model-based and model-free. Model-based methods rely on vehicle dynamics models, while model-free methods leverage data-driven techniques like deep learning. This paper focuses only on the model-free approach. Bojarski et al. [10] introduced an end-to-end steering estimation using CNN with a single RGB camera input. Capito et al. [13] improved the navigation capability by incorporating optical flow into the input. The importance of temporal information has also been highlighted in many works. Indeed, Eraqi et al. [24] employed LSTM networks to enhance steering control, while Xu et al. [25] combined fully convolutional networks with LSTM and semantic segmentation for improved road condition interpretation. Lechner et al. [4] proposed Neural Circuit Policies (NCPs), which provide interpretable decision maps from high-dimensional inputs. These studies indicate that integrating temporal information between frames can improve steering accuracy.
II-E Multi-Task Learning in Computer Vision
Recent works have explored multi-task learning (MTL) frameworks to simultaneously address multiple vision tasks, leveraging shared representations for improved performance. For instance, online knowledge distillation techniques have been proposed to mitigate negative transfer between tasks such as semantic segmentation and depth estimation, enhancing overall learning stability [26]. Joint-confidence-guided MTL frameworks have also been introduced for 3D reconstruction, combining depth prediction, semantic labeling, and surface normal estimation to improve feature fusion [27]. Moreover, collaborative MTL approaches have demonstrated effectiveness in handling object detection, segmentation, and tracking tasks by establishing associative connections among task heads [28].
Inspired by this synergy, we take a different perspective. Instead of demonstrating how MTL can enhance the performance of one or a few task through training, we hypothesize that the shared features learned from a diverse set of navigation tasks are sufficient not only for achieving good performance on each individual task but also for supporting a related navigation task—steering estimation—which is not present during training.
@@ III Methods
Given the challenge of finding an annotated dataset for all the targeted tasks, we selected the CityScapes dataset [30] for training due to its high-quality ground truth segmentation and structured video sequences. This sequential format is particularly suitable for self-supervised training in depth, optical flow, and motion mask estimation without requiring explicit labels. However, we did not evaluate the steering prediction in CityScapes. Instead, this evaluation is performed on the MIT dataset [4] after fine-tuning a prediction head (with the frozen encoder), which provides the appropriate annotations for this task. CityScapes dataset serves as a crucial foundation for training the unified encoder, allowing it to capture rich visual and motion cues that contribute to downstream tasks (e.g., steering estimation).
III-A Panoptic, Instance, Semantic Segmentations
For these three supervised segmentation tasks, we employ the OneFormer approach. This is a task-conditioned universal image segmentation model that achieves state-of-the-art performance across all segmentation tasks with a unified model and architecture. The OneFormer framework adopts a task-conditioned joint training strategy that allows simultaneous training on all segmentation tasks using a single model, reducing resource requirements significantly.
@@ III-A Panoptic, Instance, Semantic Segmentations
III-B Depth, Pose, 3D Flow Estimation, Motion Mask
@@ IV-A Training Setup
We use the KITTI Eigen split for initial ablation (trained only with the depth and the pose decoders) and CityScapes for the remaining experiments, with image resolutions of $192\times 640$ and $192\times 512$, respectively. Images are preprocessed into triples using scripts from [18]. For CityScapes, the lower 25% of images is cropped to exclude the front car [33]. The entire CityScapes dataset is used for supervised segmentation tasks. Pretraining is conducted on a single NVIDIA RTX A5000 with a batch size of 6 (3 images for the supervised tasks, i.e., segmentations, and 3 triples for the self-supervised tasks) over four steps: (1) training the shared encoder with depth, pose, and segmentation decoders for 60,000 steps, (2) training the 3D scene flow decoder for 40,000 steps with other components frozen, (3) training the pose decoder, 3D scene flow decoder, and motion mask segmentation decoder for 40,000 steps with other components frozen, and (4) training the whole network for 250,000 steps. After pretraining, we follow the tenfold cross-validation procedure by Lechner et al. [4] to fine-tune the model on the steering angle estimation task. The values of hyperparameters are: $\lambda_{\text{cls}}=2,\lambda_{\text{bce}}=\lambda_{\text{dice}}=\lambda_{\text{c}}=5,\lambda_{\text{contrast}}=0.5,\lambda_{\text{m}}=\lambda_{\text{g}}=\beta_{1}=0.1,\beta_{2}=1e-3,\lambda_{\text{1}}=\lambda_{\text{2}}=\lambda_{\text{3}}=1$.
@@ IV-B Ablation Studies
$MD2_{sep}$ 0.115 0.193 0.877
$MD2_{sh}$ 0.125 0.201 0.857
$\text{Our}_{sep}$ 0.108 0.183 0.884
$\text{Our}_{sh}$ 0.117 0.190 0.872
$\text{Our}_{sh}$ 0.109 0.184 0.884
$\text{Our}_{sep}$ 0.103 0.157 0.885
$\text{Our}_{sh}$ 0.126 0.182 0.850
$\text{Our}_{sh}$ 0.134 0.183 0.833
$\text{Our}_{sh}$ 0.106 0.158 0.888
TABLE I: 1) (Upper part) Ablation on multi-scale pose decoder using KITTI dataset. 2) (Lower part) Analysis on knowledge distilation on CityScapes. MS, FM, and KD refers multi-scale pose decoder, 3D scene Flow & Motion mask, and Knowledge Distilation. sep and sh refer to the separate and shared version. For the separate encoder, a ResNet18 was employed similar to [19].
@@
Method IM Sem. #$\#f$ D Error metric ($\downarrow$) Accuracy metric ($\uparrow$)
$Abs_{rel}$ $Sq_{rel}$ $RMSE$ $RMSE_{log}$ $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
@@
TABLE III: Depth evaluation on the KITTI (K), and CityScapes (CS) Dataset. IM and Sem. stand for independent motion and semantics. #$\#f$ indicates the number of frames during inference. D is the used datatset for training and evaluation. Bold is best. Underline is ours.
Knowledge distillation for 3D flow & motion mask. We evaluate the effectiveness of knowledge distillation in training our unified encoder on the CityScapes dataset, which includes many dynamic objects, making 3D scene flow and motion mask networks crucial. Table I-2 shows the depth error and accuracy for different models. First, using a shared encoder with our multi-scale pose decoder achieves good performance compared to using separate encoders for different tasks. Second, training the shared encoder without knowledge distillation leads to suboptimal performance, even worse than the shared encoder model that doesn’t consider 3D scene flow and motion masks. However, incorporating knowledge distillation improves the shared encoder’s performance, making its performance inline with models with separate encoders. This highlights the effectiveness of our strategy.
@@ IV-C Comparison with State-of-the-Art Methods
CNN-GRU 64 units [4] $1.25\pm 1.02$ $5.06\pm 6.64$
CNN-LSTM 64 units [4] $\mathbf{0.19}\pm 0.05$ $\mathbf{3.17}\pm 3.85$
VAE-LSTM 64 units [7] $0.54\pm 0.26$ $4.70\pm 4.80$
VAE-LSTM 19 units [7] $0.60\pm 0.30$ $6.75\pm 8.33$
(1) Swin-AttnPool (ImageNet pretrained) $2.91\pm 2.23$ $11.03\pm 10.03$
(2) Swin-AttnPool (Encoder unfrozen) $9.62\pm 3.24$ $16.46\pm 11.13$
(3) Swin-AttnPool (Encoder frozen) $1.64\pm 1.63$ $5.41\pm 6.06$
@@
IV-D Dense Latent Space to Steering: Evaluation
Table IV summarizes the performance of various models for steering prediction, including our proposed Swin-AttnPool approach and several existing CNN-based [4] and VAE-based models [7]. To evaluate the effectiveness of the features learned through our training strategy, we present three variants of the same encoder architecture:
  • ImageNet-pretrained (frozen): The encoder is pretrained on ImageNet, frozen, and fine-tuned on the steering estimation task.
  • Our pretrained (unfrozen): The encoder is initialized with our pretrained weights and unfrozen during fine-tuning.
  • Our pretrained (frozen): The encoder is initialized with our pretrained weights and remains frozen during fine-tuning.
The results reveal several key insights. First, our frozen encoder (variant 3) outperforms the frozen encoder pretrained on ImageNet (variant 1), indicating that the visual features learned through our multi-task training are more relevant for navigation tasks. Second, the frozen encoder (variant 3) also surpasses its unfrozen counterpart (variant 2), suggesting that exclusively fine-tuning for the steering task may overlook valuable features learned from other navigation-related tasks, which can positively contribute to steering performance.
Finally, although Table IV shows that our training error (variant 3) is relatively high, the test error remains competitive compared to the average performance of all methods. Indeed, our frozen encoder outperforms the VAE-LSTM (19 units) model and performs comparably to the CNN-GRU (64 units) approach. The relatively high error could be due to architectural aspects, either within the encoder or the prediction head, that may not be fully optimized for the steering estimation task.
V Discussion
Training a unified encoder for multiple tasks introduces significant optimization challenges, particularly in balancing supervised and self-supervised losses. One notable issue was the dominance of segmentation loss over motion-based tasks, resulting in imbalanced feature learning. To address this, we introduced a knowledge distillation approach from a multi-encoder teacher network, which guided the learning of motion segmentation and scene flow, ensuring stable convergence. Additionally, the proposed multi-scale pose decoder enhanced depth estimation, mitigating inaccuracies in dynamic scenes. These improvements preserved the advantages of multi-task learning while addressing convergence issues.
Our approach represents one of the first attempts to integrate human-like perceptual information into a shared encoder, achieving performance comparable to multi-encoder methods in the literature. We argue that accurate autonomous driving requires human-like perception and decision-making. By leveraging a dense feature space enriched with human-understandable representations, the model captures richer, more contextual information about the driving scene. This leads to more robust and interpretable decisions in complex real-world scenarios.
Although the strategy shows promising results, higher numerical errors were observed in certain evaluations. These discrepancies could stem from two factors: (1) the used encoder architecture may not be fully optimized for the steering estimation task, and (2) the prediction head was not extensively analyzed, as we only employed an attention pooler among many possible alternatives. Future work should explore various prediction head architectures and refine the model to align more closely with task-specific requirements, potentially reducing these errors while preserving the benefits of multi-task learning.
Additionally, while our approach demonstrates the potential of multi-task learning, it introduces increased computational complexity due to the simultaneous training of multiple tasks. Future research should explore more efficient training strategies to optimize performance and scalability.
VI Conclusion
In this work, we presented a unified encoder architecture for autonomous driving, achieved through multi-task learning that incorporates human-like visual perception necessary for navigation. By integrating knowledge from critical computer vision tasks—including depth, pose, 3D scene flow estimation, and various segmentation tasks—into a single encoder, our approach enables efficient and compact multi-task inference.